EMISSION SPECTRA FROM Cs-He EXCIPLEXES IN A COLD HELIUM GAS

\underline{K. ENOMOTO}, K. HIRANO, M. KUMAKURA, and T. YABUZAKI, Department of Physics, Kyoto University, Kyoto 606-8502, Japan.

We have observed the broadband emission spectra over the range of 9500 cm^{−1} ∼ 11750 cm^{−1} from Cs^{∗}He_{n} (n = 1, 2) exciplexes in a cold helium gas (1.3 K ∼ 100 K). The profile of the emission spectra drastically changes with varying the temperature and the helium gas density, which is due to the change of the relative population of vibrational states of Cs^{∗}He excimers. The D_{2} excitation (6 ^{2}S_{1/2} → 6 ^{2}P_{3/2}) leads to clear multi-component broadband florescence from Cs^{∗}He(A ^{2}Π_{1/2}, A ^{2}Π_{3/2}) and Cs^{∗}He_{2}. The D_{1} excitation (6 ^{2}S_{1/2} → 6 ^{2}P_{1/2}) leads to very weak broadband fluorescence from Cs^{∗}He(A ^{2}Π_{1/2}). The weakness in the case of the D_{1} excitation comes from the potential barrier of the A ^{2}Π_{1/2} state of Cs^{∗}He. We will discuss the assignment of the spectral components, the population distribution of the vibrational states of Cs^{∗}He, and the predissociation rate of the A ^{2}Π_{1/2} state, and will compare the obtained results with the theoretical calculation using the Cs-He interatomic potentials of Ref. [J. Pascale, Phys. Rev. A 28, 632 (1983)] incorporated with the spin-orbit interaction.